OpenStudy (anonymous):

Which sets of numbers are closed under subtraction? Choose all answers that are correct. A. odd natural numbers B. rational numbers C. {0, 1} D. {0, 1, 2}

zepdrix (zepdrix):

Take a sample from each set and see if you're able to leave the set :)

zepdrix (zepdrix):

Odd Naturals: 3 - 5 = -2 Oops, we fell out of the set of Odd Naturals! :O No closure. ya?

zepdrix (zepdrix):

C and D are kind of tricky. If your set contains only 0 and 1, is there any way you can leave that set by using subtraction? \(\Large\rm 1-0=0\) <-- Hmm this is ok. \(\Large\rm 0-1=-1\) <-- Oops! We left the set {0,1} :O

OpenStudy (agl202):

i think only b is correct because when you subtract two rational numbers, you always get back a rational number.